# An Adaptive Computational Intelligence Approach to Personalised Health and Immune Age Characterisation from Common Hematological Markers

### v5

@@ 6 Conclusions
We have introduced risk-assessment scores and studied their behaviour against typical synthetic and empirical disease cases. We have introduced a learning procedure to adapt normal versus abnormal reference values to personalised ranges for more precise and individual medical assessment and demonstrated the capabilities and applications of the scores for triaging purposes against a popular public database showing how the scores separate healthy from unhealthy cohorts both self-reported with significant noise and determined by typical normal reference values. The approaches were shown to be informative for quick and entry-level patient health sorting and triaging.
The numerical score was shown to be able to define an ‘immune age’, with some differences for males and females who are known to show greater fluctuations due to e.g. menopause. In the case of estimation of immune age, the chronological age of healthy people was closely predicted by common hematological markers typically measured in FBC or CBC, and for unhealthy self-reported or by typical abnormal reference values, the score was not predictive of chronological age, conforming with the theoretical expectation of a good predictor of health by the divergence between estimated immune and actual chronological age.
Unlike statistical or traditional machine learning and black-box approaches, the approach introduced is transparent to clinicians, health professionals, and health consumers who can follow the calculations and interpretation to be assessed in a clinical context for rapid screening and decision support.
@@
As this research did not involve human subjects and only publicly available data, it was exempt from an institutional review board or ethical approvals.
Funding
Santiago Hérnandez-Orozco was supported by grant SECTEI/137/2019 awarded by Subsecretaría de Ciencia, Tecnología e Innovación de la Ciudad de México (SECTEI).

### v7

@@ 5 Conclusions
We have introduced risk-assessment scores as clinical informatics and studied their behaviour against typical synthetic and empirical disease cases. We have introduced a learning procedure to adapt normal versus abnormal reference values to personalised ranges for more precise and individual medical assessment, and demonstrated the capabilities and digital applications of the scores for triaging purposes against a popular public database, showing how the scores separate healthy from unhealthy cohorts, both as self-reported, with significant noise, and as determined by typical normal reference values. The approaches were shown to be informative for quick and entry-level patient health sorting and triaging.
As we have shown, the analyte-driven score performance underscores the potential of these commonly tested blood analytes as reliable biomarkers for health status prediction, providing critical insights for early detection and precision medicine.
Our findings suggest that when combined with feed-forward ANN and Transformer models, these are well-suited for leveraging the NIS immune score to predict health status. The performance of these deep learning models can be attributed to their ability to capture complex, non-linear relationships within the training data, which simpler models cannot detect. This highlights the potential of integrating advanced hybrid neuro-symbolic AI models guided by closed, explainable and controllable algorithms, such as the risk assessment scores we have introduced.
The numerical score was also shown to be able to define an ‘immune age’ as a marker for health according to the classification test, with differences expected and reported for males and females, who are known to show greater fluctuations due to, e.g., menopause, for which the recommendation is to take these life events into consideration for bias-correction purposes. In the case of the estimation of immune age, the chronological age of healthy people was closely predicted by the scores based on common haematological markers typically massively measured in FBC or CBC tests. For unhealthy people, whether as self-reported or as measured by typical abnormal reference values, the score was not predictive of chronological age, conforming with the theoretical expectation of a good predictor of health by divergence from chronological age in the domain of digital medicine.
Unlike statistical or traditional machine learning and black-box approaches, the hybrid approach introduced here helps guide the search, is more transparent to and controllable by health professionals or automated supervisors and by health consumers, facilitating rapid screening, regular and remote patient monitoring and decision support.
@@
As this research did not involve human subjects and used only publicly available anonymized data, it was exempt from consideration by an institutional review board and from ethical approvals.